If you invest $10,000 and it grows at annual rate of 15% (compounded annually), how many months it will take to grow to $20,000?
Answer #1

future value of an amount = present value *(1+interest rate)^n

here,

future value = $20,000

present value =$10,000

interest rate = 15% per annum

=>1.25% per month.

=>0.0125

n = number of periods = to be found out.

now,

$20,000 =$10,000*(1.0125)^n

=>2 = (1.0125)^n

=>log 2 =n log (1.0125)

=>0.30103= n (0.005395032)

=>n = 0.30103/0.005395032

=>n = 55.7976301

=>55.80 months....(rounded to two decimals).